# ADSeeker: A Knowledge-Infused Framework for Anomaly Detection and Reasoning

### v1

@@ 4.4 Hierarchical Sparse Prompt
CLIP 74.1 66.4 34.5 54.3 73.9 56.5 78.4 62.6
WinCLIP 91.8 78.8 68.2 63.6 92.6 90.0 80.5 80.8
AnomalyCLIP 91.5 82.1 88.3 77.0 90.3 93.4 94.6 88.2
AdaCLIP 89.2 85.8 88.6 76.0 94.8 91.4 97.7 89.1
Anomaly-OV 94.0 91.1 89.0 81.7 93.9 97.6 95.5 91.8
@@
As illustrated in Figure 3 (B), given the query image $I_{Q}$, the defect features of query image are passed through linear layers to obtain the query embeddings $E_{q}\in{R}^{B\times d}$. Then we introduce learnable prompt embeddings $E_{l}\in{R}^{K\times d}$ and adaptive parameters $P_{i}\in{R}^{B\times K}$ which are then fed into the sparse optimization module. The prompt embeddings $E_{l}$ extract essential features from the embeddings $E_{q}$ through $\mathcal{N}$ iterative updates. At the $n_{th}$ round of iteration, the current residual $R_{n}$ can be obtained by:
@@
The final embeddings $E_{f}$ are integrated to a certain depth during the forward pass in the vision and text encoder. As we explained in the section 4.1, the type-level textual prompts will be concatenated with defect-region features to enhance ADSeeker’s fine-grained understanding of the anomaly pattern. Extensive experiments have demonstrated the potential of this module in bridging the gap between human anomaly inspector and ADSeeker.
@@ 5.1 Zero-shot Anomaly Detection
This study compares the proposed ADSeeker with existing ZSAD approaches. We conduct extensive experiments on publicly available AD benchmarks from industrial and medical domains. Specifically, for the industry domain, we utilize MVTec AD, VisA, BTAD, and MPDD datasets. In the medical domain, we re-masked the brain medical detection datasets BrainMRI, HeadCT, and Br35H, making them qualified anomaly detection datasets. The anomaly detection and localization performance of ADSeeker is evaluated by the Area Under the Receiver Operating Characteristic Curve (AUROC). The experiment follows the Zero-shot setting, our ADSeeker is fine-tuned on MVTec AD to evaluate the ZSAD performance on other datasets. When evaluating the model on MVTec AD, the corresponding train-set is replaced by VisA. To ensure the experiment fairness, we stopped utilizing Q2K RAG to ensure test-set is not leaked during test.
The results are presented in Table 2, our ADSeeker have achieved significant enhancement on most of the AD benchmarks in both medical and industrial domains. The enhancement of our model mainly originates from the Hierarchical Sparse Prompt mechanism, which enables the model to extract the defect-region features and learn more fine-grained semantics for anomaly detection in interaction with learnable textual prompts. ADSeeker has emerged as the top performer, boasting the highest average ranking. This achievement highlights its impressive ability to generalize effectively across various domains.
@@ 5.2 Anomaly Reasoning & ADSeeker Framework
Table 3: Anomaly reasoning performance comparison with and the best-performing training methods. * represents base model under the Seek-Setting
ADSeeker LoRA Accuracy
Baseline Q2K RAG Expert
@@
Table 4: Ablation Results of proposed modules and training settings. The ✓ in each columns signifies module integration. The number in “LoRA” column denotes we utilize LoRA to fine-tune LLM for that epochs.
Model Scale Setting Anomaly Defect Object Average
LLaVA-OV 7B Vanilla 51.17 46.13 41.85 62.19 69.73 90.31 80.93 63.19
LLaVA-NeXT 7B Vanilla 57.64 33.79 47.72 51.84 67.93 81.39 74.91 59.32
InternVL2 8B Vanilla 59.97 43.85 47.91 57.60 78.10 74.18 80.37 63.14
Qwen2.5-VL 3B Vanilla 51.10 41.07 44.87 61.43 75.99 86.54 79.58 62.94
ADSeeker Seek-Setting 58.17 (+7.07) 48.71 (+7.64) 53.67 (+8.80) 69.79 (+8.36) 80.36 (+4.37) 86.78 (+0.24) 82.25 (+2.67) 68.53(+5.59)
Qwen2.5-VL 7B Vanilla 55.35 49.86 54.34 57.18 73.20 92.07 84.31 66.62
ADSeeker Seek-Setting 68.62 (+13.27) 53.82 (+3.96) 55.57 (+1.23) 60.51 (+3.33) 77.40 (+4.20) 89.91 (-2.16) 83.48 (-0.83) 69.90(+3.28)
Table 5: Performance comparison of both ADSeeker and other open-source MLLMs in MMAD with the standard 1-shot setting. Notably, each model is also tested under Seek-Setting.
With the powerful capability of retrieving relevant expertise and the anomaly prior feature from the AD Expert module, ADSeeker has achieved great improvement in performing anomaly reasoning tasks. As is shown in Figure 2, ADSeeker could accurately locate defects and point out the defect type. Its anomaly description could achieve the accuracy of human anomaly inspectors.
We evaluated anomaly reasoning capability via accuracy on multiple-choice subtasks from the MMAD benchmark, comprising four anomaly-related and two object-related subtasks. Under a default 1-shot setting, compared models received a randomly selected normal sample as template alongside the query image. For ADSeeker and Seek-Setting models, the retrieved knowledge document replaced this normal sample as the template. As shown in Table 5, ADSeeker (i.e., Qwen2.5-VL under Seek-Setting) was compared against open-source MLLMs. And each model is also tested under Seek-Setting equipped with our Q2K RAG and AD Expert module to validate the efficacy of the plug-and-play framework. ADSeeker delivers a remarkable accuracy of 69.90%, excelling across various subtasks. And our framework shows enhance mainly on defect classification, localization and Anomaly Discrimination. The injected anomaly prior feature enhances defect localization through visual information fusion, while the retrieved knowledge document—containing specific defect types—facilitates defect classification and anomaly discrimination.
Pretraining remains the most common solution for anomaly reasoning. Recent best-performing training-based approaches (e.g., Anomaly-OV and Anomaly-R1) have achieved remarkable results in IAD tasks. To evaluate the efficacy of our framework, we compare their anomaly reasoning performance under both original training strategies and the Seek-Setting. Due to the limitation of SEEK-M&V content, we utilize part of the MMAD dataset to conduct evaluation..As illustrated in Table 3, extensive experiments demonstrate that ADSeeker could outperform existing SOTA IAD approaches without additional training. Notably, Anomaly-R1 achieves remarkable performance by adapting the GRPO strategy to fine-tuning with a limited training set, the utilization of that strategy is suggested for future works.
5.3 Efficiency and Efficacy Analysis
To demonstrate the efficacy of each proposed module in ADSeeker, we conduct ablation experiments by comparing the accuracy of ADSeeker in 2 subtasks among MMAD. We primarily focus on 3 aspects: the Q2K RAG, the AD Expert module and the impact of LoRA fine-tuning on ADSeeker. Our team constructed a 43K small-scale instruction tuning dataset, we collected domain knowledge from AD datasets and our SEEK-M&V content. The dataset content contains most of the knowledge base content and is constructed using template examples. As shown in Table 4, the full-configuration ADSeeker achieves optimal performance across tasks. Each module contributes uniquely to anomaly reasoning: Q2K RAG retrieves domain knowledge, while the AD Expert module provides anomaly prior features for localization and discrimination. Note that model capability peaks then declines with extended training epochs. And small-scale training sets exacerbate overfitting and catastrophic forgetting of pretrained knowledge. Moreover, our Q2K RAG framework surpasses LoRA fine-tuning in both generalizability and precision.
As is shown in Table 6, we calculate the Memory Usage and Inference Time of each module to demonstrate the efficiency of our proposed framework. When deployed on LLMs of comparable scale, ADSeeker incurs modest overhead: $\leq 27\%$ additional memory consumption and $\leq 2s$ average inference latency. This demonstrates strong potential in real-time industrial applications. Furthermore, ADSeeker is a training-efficient framework, which shows no need for large-scale pre-training and effectively reduces the consumption of computational resources. In general, proposed as a plug-and-play framework, ADSeeker could achieve the trade-offs between performance and computational resources.
@@ 5.3 Efficiency and Efficacy Analysis
Table 6: Memory usage and inference time calculation of each module (mean±SD)
@@ 5.4 Influence of Type-level Feature
We conduct comprehensive evaluations to validate the efficacy of the type-level representation. On the left of the Figure 5, we calculate the similarity of the patch tokens and different kinds of textual prompts. Experiments reveal stronger alignment between normal samples and [obj]-only prompts, while anomalous samples exhibit peak similarity to prompts containing both [cls]&[obj]. Furthermore, for defect samples, [cls] prompts demonstrate significantly higher feature value than [obj] prompts. These findings empirically validate the efficacy of type-level information representation. Our cross-dataset experiments of different-level prompts efficacy reveals that type-level [cls] features significantly outperform [obj] features in image-level anomaly detection tasks. However, [obj] features maintain a noteworthy complementary role, as evidenced by the ablation study in Figure 5.
Figure 5: Visualization the efficacy of the type-level feature through the Similarity Matrix and Ablation Study
@@ 6 Conclusion
In this paper, we propose an anomaly task assistant ADSeeker, leveraging the multimodal Q2K RAG to seek for domain knowledge. To equip ADSeeker with AD expertise, we establish the first visual document knowledge base SEEK-M&V for MLLMs in IAD tasks. Additionally, the HSP module is proposed to extract type-level features and generate sparse prompts to enhance the ZSAD performance. Our work delves into the potential application of multimodal RAG in anomaly detection, introducing a knowledge-infused framework without additional training.

### v2

@@ 4.4 Hierarchical Sparse Prompt
CLIP[29] 74.1 66.4 34.5 54.3 73.9 56.5 78.4 62.6
WinCLIP[15] 91.8 78.8 68.2 63.6 92.6 90.0 80.5 80.8
AnomalyCLIP[40] 91.5 82.1 88.3 77.0 90.3 93.4 94.6 88.2
AdaCLIP[6] 89.2 85.8 88.6 76.0 94.8 91.4 97.7 89.1
AnomalyOV[34] 94.0 91.1 89.0 81.7 93.9 97.6 95.5 91.8
@@
As illustrated in Fig. 4(B), given the query image $I_{Q}$, the defect features of query image are passed through linear layers to obtain the query embeddings $E_{q}\in{R}^{B\times d}$. Then we introduce learnable prompt embeddings $E_{l}\in{R}^{K\times d}$ and adaptive parameters $P_{i}\in{R}^{B\times K}$ which are then fed into the sparse optimization module. The prompt embeddings $E_{l}$ extract essential features from the embeddings $E_{q}$ through $\mathcal{N}$ iterative updates. At the $n_{th}$ round of iteration, the current residual $R_{n}$ can be obtained by:
@@
The final embeddings $E_{f}$ are integrated to a certain depth during the forward pass in the vision and text encoder. As we explained in the Sec. 5.1, the type-level textual prompts will be concatenated with defect-region features to enhance ADSeeker’s fine-grained understanding of the anomaly pattern. Extensive experiments have demonstrated the potential of this module in bridging the gap between human anomaly inspector and ADSeeker.
Model Scale Setting Anomaly Defect Object Average
LLaVA-OV[21] 7B Baseline 51.17 46.13 41.85 62.19 69.73 90.31 80.93 63.19
LLaVA-NeXT[25] 7B Baseline 57.64 33.79 47.72 51.84 67.93 81.39 74.91 59.32
InternVL2[9] 8B Baseline 59.97 43.85 47.91 57.60 78.10 74.18 80.37 63.14
Qwen2.5-VL[3] 3B Baseline 51.10 41.07 44.87 61.43 75.99 86.54 79.58 62.94
Seek-Setting 58.17 (+7.07) 48.71 (+7.64) 53.67 (+8.80) 69.79 (+8.36) 80.36 (+4.37) 86.78 (+0.24) 82.25 (+2.67) 68.53(+5.59)
Qwen2.5-VL[3] 7B Baseline 55.35 49.86 54.34 57.18 73.20 92.07 84.31 66.62
Seek-Setting 68.62 (+13.27) 53.82 (+3.96) 55.57 (+1.23) 60.51 (+3.33) 77.40 (+4.20) 89.91 (-2.16) 83.48 (-0.83) 69.90(+3.28)
Table 2: Anomaly Reasoning Performance. Performance comparison of both ADSeeker (Qwen2.5-VL under Seek-Setting) and other open-source MLLMs in MMAD with the standard 1-shot setting. Notably, each model is also tested under Seek-Setting.
@@ 5.1 Zero-shot Anomaly Detection
We conduct extensive experiments on publicly available AD benchmarks from industrial and medical domains. Specifically, for the industry domain, we utilize MVTec AD[4], VisA[41], BTAD[26], and MPDD[16] datasets. In the medical domain, we re-masked the brain medical detection datasets BrainMRI[2], HeadCT[1], and Br35H[13], making them qualified anomaly detection datasets. The experiment follows the Zero-shot setting, our ADSeeker is fine-tuned on MVTec AD to evaluate the ZSAD performance on other datasets. When evaluating the model on MVTec AD, the corresponding train-set is replaced by VisA. To ensure the experiment fairness, we stopped utilizing Q2K RAG to ensure test-set is not leaked during test. We will demonstrate our complete experimental setup and training parameters.
The results are presented in Tab. 1, our ADSeeker have achieved significant enhancement on most of the AD benchmarks in both medical and industrial domains. The enhancement of our model mainly originates from the Hierarchical Sparse Prompt mechanism, which enables the model to extract the defect-region features and learn more fine-grained semantics for anomaly detection in interaction with learnable textual prompts. ADSeeker has emerged as the top performer, boasting the highest average ranking. This achievement highlights its impressive ability to generalize effectively across various domains, as the visualization results shown in Fig. 6.
@@ 5.2 Anomaly Reasoning & ADSeeker Framework
Table 3: Comparison with training-methods. Anomaly reasoning performance comparison with and the best-performing training methods. * represents base model under the Seek-Setting.
With the powerful capability of retrieving relevant expertise and the anomaly prior feature from the AD Expert module, ADSeeker has achieved great improvement in performing anomaly reasoning tasks. As is shown in Fig. 2, ADSeeker could accurately locate defects and point out the defect type. Its anomaly description could achieve the accuracy of human anomaly inspectors.
We evaluated anomaly reasoning capability via accuracy on multiple-choice subtasks from the MMAD benchmark, comprising four anomaly-related and two object-related subtasks. Under a default 1-shot setting, compared models received a randomly selected normal sample as template alongside the query image. For ADSeeker and Seek-Setting models, the retrieved knowledge document replaced this normal sample as the template. As shown in Tab. 2, ADSeeker (i.e., Qwen2.5-VL under Seek-Setting) was compared against open-source MLLMs. And each model is also tested under Seek-Setting equipped with our Q2K RAG and AD Expert module to validate the efficacy of the plug-and-play framework. ADSeeker delivers a remarkable accuracy of 69.90%, excelling across various subtasks. And our framework shows enhance mainly on defect classification, localization and Anomaly Discrimination. The injected anomaly prior feature enhances defect localization through visual information fusion, while the retrieved knowledge document—containing specific defect types—facilitates defect classification and anomaly discrimination.
Pretraining remains the most common solution for anomaly reasoning. Recent best-performing training-based approaches (e.g., Anomaly-OV and Anomaly-R1) have achieved remarkable results in IAD tasks. To evaluate the efficacy of our framework, we compare their anomaly reasoning performance under both original training strategies and the Seek-Setting. Due to the limitation of SEEK-M&V content, we utilize part of the MMAD dataset to conduct evaluation. As illustrated in Tab. 3, extensive experiments demonstrate that ADSeeker could outperform existing SOTA IAD approaches without additional training. Notably, Anomaly-R1 achieves remarkable performance by adapting the GRPO strategy to fine-tuning with a limited training set, the utilization of that strategy is suggested for future works.
5.3 Efficiency and Efficacy Analysis
ADSeeker LoRA[14] Accuracy
Baseline[3] Q2K RAG Expert
@@ 5.3 Efficiency and Efficacy Analysis
Table 4: Ablation Results of proposed modules and training settings. The ✓ in each columns signifies module integration. The number in “LoRA” column denotes we utilize LoRA to fine-tune MLLM for that epochs.
@@
Table 5: Quantification of reasoning efficiency. Memory usage and inference time calculation of each module (mean±SD).
To demonstrate the efficacy of each proposed module in ADSeeker, we conduct ablation experiments by comparing the accuracy of ADSeeker in 2 subtasks among MMAD. We primarily focus on 3 aspects: the Q2K RAG, the AD Expert module and the impact of LoRA fine-tuning on ADSeeker. Our team constructed a 43K small-scale instruction tuning dataset, we collected domain knowledge from AD datasets and our SEEK-M&V content. The dataset content contains most of the knowledge base content and is constructed using template examples. As shown in Tab. 4, the full-configuration ADSeeker achieves optimal performance across tasks. Each module contributes uniquely to anomaly reasoning: Q2K RAG retrieves domain knowledge, while the AD Expert module provides anomaly prior features for localization and discrimination. Note that model capability peaks then declines with extended training epochs. And small-scale training sets exacerbate overfitting and catastrophic forgetting of pretrained knowledge. Moreover, our Q2K RAG framework surpasses LoRA fine-tuning in both generalizability and precision.
As is shown in Tab. 5, we calculate the Memory Usage and Inference Time of each module to demonstrate the efficiency of our proposed framework. When deployed on LLMs of comparable scale, ADSeeker incurs modest overhead: $\leq 27\%$ additional memory consumption and $\leq 2s$ average inference latency. This demonstrates strong potential in real-time industrial applications. Furthermore, ADSeeker is a training-efficient framework, which shows no need for large-scale pre-training and effectively reduces the consumption of computational resources. In general, proposed as a plug-and-play framework, ADSeeker could achieve the trade-offs between performance and computational resources.
Figure 6: Validation of type-level feature representation.
@@ 5.4 Influence of Type-level Feature
On the left of the Fig. 6(A), we calculate the similarity of the patch tokens and different kinds of textual prompts. Experiments reveal stronger alignment between normal samples and [obj]-only prompts, while anomalous samples exhibit peak similarity to prompts containing both [cls]&[obj]. Furthermore, for defect samples, [cls] prompts demonstrate significantly higher feature value than [obj] prompts. These findings empirically validate the efficacy of type-level information representation. Our cross-dataset experiments of different-level prompts efficacy reveals that type-level [cls] features significantly outperform [obj] features in image-level anomaly detection tasks. However, [obj] features maintain a noteworthy complementary role, as evidenced by the ablation study in Fig. 6.
As shown in Fig. 6(B), we utilize the type-level feature to detect anomaly industry images and provide visualization results. Combining the Image-Level AUROC in Tab. 1 and visualization results, we concluded that the ability to extract general features is the key to achieving zero-shot rendering.
@@ 6 Conclusion
In this paper, we propose an anomaly task assistant ADSeeker, leveraging the visual document Q2K RAG to seek for domain knowledge. To equip ADSeeker with expertise, we establish the first visual document knowledge base SEEK-M&V for MLLMs in IAD tasks. Our work delves into the potential application of external knowledge in anomaly reasoning and detection, introducing a knowledge-grounded reasoning framework without large-scale training.
7 Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grant No. U24A20221), the Key R&D Program of Shandong Province, China (Grant No. 2023CXGC010112), the Distinguished Young Scholar of Shandong Province (Grant No. ZR2023JQ025), the Taishan Scholars Program (Grant No. tstp20250708), and the Major Basic Research Projects of Shandong Province (Grant No. ZR2022ZD32).